# GP-MoLFormer-Sim: Test Time Molecular Optimization through Contextual Similarity Guidance

### v1

@@ 2.1 Background information — GP-MoLFormer
2.2 Target-guided generation with GP-MoLFormerGP-MoLFormer-Sim
@@
Note that the cosine similarity is used to “tilt” the logits of the GP-MoLFormer using similarity to a neighborhood formed by target molecules in the embedding space. We can push this idea further using a kernel density estimator (KDE) with a gaussian kernel. The temperature $T$ of the kernel induces further locality control. We can approximate the KDE using Random Fourier Features (rff, ) (lines 8 and 14 in Algorithm 1).
@@ 4.1 Similarity-guided molecule generation
Table 1 reports the mean similarity ($T_{sim}$) of top-k most similar generations obtained using GP-MoLFormer-Sim (GPMFS) and baseline methods. Similarity values of the top-ranked generations show that target similarity-guided decoding using GPMFS performs better than test-time baselines like random sampling and random search. The proposed method also outperforms a graph generative model that is RL-tuned to optimize the target similarity (S model) across all values of k up to $10^{4}$. The reported min, mean, and max QED values show the inverse relation between QED and similarity for these targets, given the mean QED value of these five targets is only 0.234. Nevertheless, 132 molecules are found to have a QED value $>0.7$ in the top-10000 most similar molecules generated by GPMFS. Those show a mean similarity of 0.47 and a max similarity of 0.67, showcasing the potential of the proposed method to guide generations on-the-fly toward a target molecule, while yielding useful molecules. A small sample of molecules generated by the GPMFS and their respective targets are visualized in Figure 2 for varying parameter settings (final parameter values can be found in the Appendix Table 16).
@@ 4.2 Sample-efficient molecular optimization — PMO benchmark
Table 2 reports the performance of GPMFS+GA on the 23 tasks from the PMO benchmark. Since the proposed method relies on a combination of the test-time steering of the deep generative model and a modified genetic algorithm, in the main article we show comparison of the proposed method with GA-based baselines specifically designed for molecular design that do not require any training of the generative model. The baselines shown in Table 2 are Graph GA jensen2019graph , STONED Nigam_Pollice_Krenn_Gomes_Aspuru-Guzik_2021 , SynNet gao2021amortized , Mol-GA tripp2023geneticalgorithmsstrongbaselines , and MOLLEO wang2024efficient . We report the rank per task based on the top-10 AUC score obtained with a maximum of 10k oracle calls for each method. Average rank and average score over all tasks are also reported in Table 2. For comparison with additional baselines and more analyses on the optimization runs, see Appendix. Results show that the proposed method scores second among GA baselines, while MOLLEO that uses Graph-GA with GPT-4 comes first. On three tasks, namely GSK3, JNK3, and ranolazine_mpo, GPMFS+GA outperforms all baselines, while on another 9 tasks it ranks second. Given the computational and actual dollar cost associated with calling GPT-4, GPMFS+GA appears as a more cost-effective alternative.
@@ 4.3 Comparison with MOLLEO in the black-box oracle setting
MOLLEO includes natural language prompting of the LLM to generate proposals based on the GA operations — crossover and mutation. While doing so, the prompt includes information about the task and the oracle function. For example, for the thiothixene rediscovery task, the prompt used in reference wang2024efficient includes the following (emphasis ours): “OBJECTIVE: has a higher thiothixene rediscovery score. TASK: thiothixene rediscovery scores. OBJECTIVE_DEFINITION: The thiothixene rediscovery score measures a molecule’s Tanimoto similarity with thiothixene’s SMILES to check whether it could be rediscovered.” This contextual information present in the prompt weakens the black-box nature of the oracle used in the PMO benchmark as the GPT-4 model has memorized the thiothixene SMILES, which is otherwise never disclosed to the other baselines. Therefore, to enable a fair comparison, we revise the prompt such that the name of the target molecule is redacted from the prompt. Results are reported in Table 3 for two exemplar tasks, namely thiothixene rediscovery and mestranol similarity. On both tasks, MOLLEO’s performance (using a gpt-4.1-mini model) drops by $\sim$33and becomes worse compared to GPMFS+GA when the molecule name is redacted. This result implies that MOLLEO’s performance depends on the LLM’s utilization of the task-relevant contextual information for proposing offspring. In contrast, our proposed method only uses the score from the (black-box) oracle, consistent with the setting of the PMO benchmark, to produce candidates during optimization and outperforms MOLLEO in that mode.
@@ 4.4 Ablation experiments
Table 5: Overall improvement due to specific method combinations in terms of average 1-based rank and top-10 AUC metric in GP-MoLFormer-Sim. Includes adding “XO” as crossover, “RFF768” as 768-dimensional Random Fourier Features, and “DIV” as diversity augmentation to the vanilla Guided Generation (GG).
@@ 6 Conclusions
Appendix A Extended Results — PMO Tasks
A.1 Optimization trajectory — an example
@@
A.2 Redacted MOLLEO Prompts
@@ A.3.1 Optimization curves
A.3.2 Extended baseline comparison
@@
Table 8 shows a comparison of our GPMFS-GA to [31]. Tables 9 through 12 show results reported in [6]. Tables 13-14 compare results from [16], and Table 15 compares results published in [33].
Table 8: Comparison of GP-MoLFormer-Sim+GA to main results in [31]. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 9: Comparison of GP-MoLFormer-Sim+GA to main results in [6] - Page 1 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 10: Comparison of GP-MoLFormer-Sim+GA to main results in [6] - Page 2 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 11: Comparison of GP-MoLFormer-Sim+GA to main results in [6] - Page 3 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 12: Comparison of GP-MoLFormer-Sim+GA to main results in [6] - Page 4 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 13: Comparison of GP-MoLFormer-Sim+GA to main results in [16] - Page 1 of 2. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 14: Comparison of GP-MoLFormer-Sim+GA to main results in [16] - Page 2 of 2. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 15: Comparison of GP-MoLFormer-Sim+GA to main results in [33]. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
A.4 Extended Results — Trypsin inhibitors
@@
Appendix B Hyperparameter Settings
@@
GP-MoLFormer embedding dimensionality Float 768 See [28]

### v2

@@ 2.1 Background information — GP-MoLFormer
2.2 Target-guided generation with $GP-MoLFormer$-Sim
@@
Note that the cosine similarity is used to “tilt” the logits of the GP-MoLFormer using similarity to a neighborhood formed by target molecules in the embedding space. We can push this idea further using a kernel density estimator (KDE) with a gaussian kernel. The temperature $T$ of the kernel induces further locality control. We can approximate the KDE using Random Fourier Features (rff) (lines 8 and 14 in Algorithm 1).
@@
Theoretical Justification
It can be shown that the computations on lines 23–24 in Algorithm 1 represent a solution of a closed-form optimization problem involving token similarity to the target molecule distribution and a KL divergence from the smoothed GP-MoLFormer distribution. Full detail is laid out in Appendix A.
@@ 4.1 Similarity-guided molecule generation
Table 1 reports the mean similarity ($T_{sim}$) of top-k most similar generations obtained using GP-MoLFormer-Sim (GPMFS) and baseline methods. Similarity values of the top-ranked generations show that target similarity-guided decoding using GPMFS performs better than test-time baselines like random sampling and random search. The proposed method also outperforms a graph generative model that is RL-tuned to optimize the target similarity (S model) across all values of k up to $10^{4}$. The reported min, mean, and max QED values show the inverse relation between QED and similarity for these targets, given the mean QED value of these five targets is only 0.234. Nevertheless, 132 molecules are found to have a QED value >0.7>0.7 in the top-10000 most similar molecules generated by GPMFS. Those show a mean similarity of 0.47 and a max similarity of 0.67, showcasing the potential of the proposed method to guide generations on-the-fly toward a target molecule, while yielding useful molecules. A small sample of molecules generated by the GPMFS and their respective targets are visualized in Figure 2 for varying parameter settings (final parameter values can be found in the Appendix Table 17).
@@ 4.2 Sample-efficient molecular optimization — PMO benchmark
Table 2 reports the performance of GPMFS+GA on the 23 tasks from the PMO benchmark. Since the proposed method relies on a combination of the test-time steering of the deep generative model and a modified genetic algorithm, in the main article we show comparison of the proposed method with GA-based baselines specifically designed for molecular design that do not require any training of the generative model. The baselines shown in Table 2 are Graph GA jensen2019graph, STONED Nigam_Pollice_Krenn_Gomes_Aspuru-Guzik_2021, SynNet gao2021amortized, Mol-GA tripp2023geneticalgorithmsstrongbaselines, and MOLLEO wang2024efficient. We report the rank per task based on the top-10 AUC score obtained with a maximum of 10k oracle calls for each method. Average rank and average score over all tasks are also reported in Table 2. For comparison with additional baselines and more analyses on the optimization runs, see Appendix. Results show that the proposed method scores second among GA baselines, while MOLLEO that uses Graph-GA with GPT-4 comes first. On three tasks, namely GSK3, JNK3, and ranolazine_mpo, GPMFS+GA outperforms all baselines, while on another 9 tasks it ranks second. Given the computational and actual dollar cost associated with calling GPT-4, GPMFS+GA appears as a more cost-effective alternative. To gauge synthesizability, average Synthetic Accessibility scores of the top 100 generated molecules in each task are reported in Appendix Table 16.
@@ 4.3 Comparison with MOLLEO in the black-box oracle setting
MOLLEO includes natural language prompting of the LLM to generate proposals based on the GA operations — crossover and mutation. While doing so, the prompt includes information about the task and the oracle function. For example, for the thiothixene rediscovery task, the prompt used in reference wang2024efficient includes the following (emphasis ours): “OBJECTIVE: has a higher thiothixene rediscovery score. TASK: thiothixene rediscovery scores. OBJECTIVE_DEFINITION: The thiothixene rediscovery score measures a molecule’s Tanimoto similarity with thiothixene’s SMILES to check whether it could be rediscovered.” This contextual information present in the prompt weakens the black-box nature of the oracle used in the PMO benchmark as the GPT-4 model has memorized the thiothixene SMILES, which is otherwise never disclosed to the other baselines. Therefore, to enable a fair comparison, we revise the prompt such that the name of the target molecule is redacted from the prompt. Results are reported in Table 3 for two exemplar tasks, namely thiothixene rediscovery and mestranol similarity. On both tasks, MOLLEO’s performance (using a gpt-4.1-mini model) drops by $\sim$33% and becomes worse compared to GPMFS+GA when the molecule name is redacted. This result implies that MOLLEO’s performance depends on the LLM’s utilization of the task-relevant contextual information for proposing offspring. In contrast, our proposed method only uses the score from the (black-box) oracle, consistent with the setting of the PMO benchmark, to produce candidates during optimization and outperforms MOLLEO in that mode.
@@ 4.4 Ablation experiments
Table 5: Overall improvement due to specific method combinations in terms of average 1-based rank and top-10 AUC metric in GP-MoLFormer-Sim. Includes adding “XO” as crossover, “RFF768” as 768-dimensional Random Fourier Features, and “DIV” as diversity augmentation to the vanilla Guided Generation ($G$).
@@ 6 Conclusions
Appendix A Theoretical Justification for Algorithm 1
Let $V$ be the vocabulary size. Given previously generated tokens represented as context $c$, let $\pi^{\text{target}}_{\mathsf{KDE}}(\cdot|c)$ be the Kernel Density Estimator (KDE) of the target molecules defined in the embedding space of GP-MoLFormer noted below as $\pi_{\mathrm{ref}}$. Our algorithm can be seen at each token generation as the closed form solution of the following optimization problem:
$\mathsf{P:}\max_{p\in\Delta_{V}}$ $(1-\alpha)\underbrace{\sum_{i=1}^{V}p_{i}\pi^{\text{target}}_{\mathrm{KDE}}(x_{i}|c)}_{\text{Similarity to target }\textsf{KDE}}-\underbrace{\mathsf{KL}(p||\pi_{\mathrm{ref}}^{\alpha}(\cdot|c))}_{\mathsf{KL}\textsf{ to {GP-MoLFormer}}}$
$\mathsf{s.t.}~\Delta_{V}=\left\{p\in[0,1]^{V}\Big|\sum_{i=1}^{V}p_{i}=1\right\},$
where for $p\in\Delta_{V}$ and $q\in\mathbb{R}_{+}^{V},$ we define the generalized KL divergence [peyré2020computationaloptimaltransport] as:
$\mathrm{KL}(p\,\|\,q)=\sum_{i=1}^{V}p_{i}\log\left(\frac{p_{i}}{q_{i}}\right)+\sum_{i=1}^{V}q_{i}-1.$
The optimization problems $\mathsf{P}$ balances between closeness to GP-MoLFormer distribution smoothed by $\alpha$ ($\pi_{\mathrm{ref}}^{\alpha}$) and the similarity to the target molecules kernel density estimator ($\pi^{\text{target}}_{\mathrm{KDE}}$). We use the notion of generalized divergence since $\pi_{\mathrm{ref}}^{\alpha}$ is a smoothed distribution that is no longer normalized. Indeed the optimal solution of this problem [ko2025large] is proportional to $\pi_{\mathrm{ref}}^{\alpha}\exp((1-\alpha)\pi^{\text{target}}_{\mathrm{KDE}})=\exp((1-\alpha)\pi^{\text{target}}_{\mathrm{KDE}}+\alpha\log\pi_{\mathrm{ref}})$. This justifies the computations on lines 23–24 in Algorithm 1, where the new guided probability is obtained using a softmax on the interpolation between the KDE estimator of the target molecules and the logits of GP-MoLFormer.
Appendix B Extended Results — PMO Tasks
B.1 Optimization trajectory — an example
@@
B.2 Redacted MOLLEO Prompts
@@ B.3.1 Optimization curves
B.3.2 Extended baseline comparison
@@
Table 8 shows a comparison of our GPMFS-GA to [tripp2023geneticalgorithmsstrongbaselines]. Tables 9 through 12 show results reported in [gao2022sample]. Tables 13-14 compare results from [kim2024geneticguidedgflownetssampleefficient], and Table 15 compares results published in [wang2024efficient].
Table 8: Comparison of GP-MoLFormer-Sim+GA to main results in [tripp2023geneticalgorithmsstrongbaselines]. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 9: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 1 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 10: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 2 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 11: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 3 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 12: Comparison of GP-MoLFormer-Sim+GA to main results in [gao2022sample] - Page 4 of 4. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 13: Comparison of GP-MoLFormer-Sim+GA to main results in [kim2024geneticguidedgflownetssampleefficient] - Page 1 of 2. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 14: Comparison of GP-MoLFormer-Sim+GA to main results in [kim2024geneticguidedgflownetssampleefficient] - Page 2 of 2. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 15: Comparison of GP-MoLFormer-Sim+GA to main results in [wang2024efficient]. Mean ($\pm$ standard deviation) AUC top-10 over 5 runs for each.
@@
Table 16: The mean ($\pm$ standard deviation) Synthetic Accessibility scores (SA) of top 100 molecules generated for each task using GP-MoLFormer-Sim+GA. The mean values are less than the threshold of 6 suggested in [Ertl_Schuffenhauer_2009], indicating that the molecules are easily synthesizable.
Task SA
albuterol 4.235 (0.433)
amlodipine 3.453 (0.101)
celecoxib 2.422 (0.322)
deco_hop 3.831 (0.541)
DRD2 4.004 (0.401)
fexofenadine 5.285 (0.320)
GSK3 3.230 (0.425)
isomers_c7 2.834 (0.795)
isomers_c9 4.301 (0.226)
JNK3 (zincseed)+RFF768+XO+DIV 3.114 (0.371)
median1 (zincseed)+RFF768+XO+DIV 3.587 (0.041)
median2 (zincseed)+RFF768+XO+DIV 2.882 (0.353)
mestranol 3.836 (0.203)
osimertimib 4.031 (0.155)
perindopril 3.791 (0.289)
QED 2.417 (0.314)
ranolazine 4.040 (0.292)
scaffold_hop 3.621 (0.273)
sitagliptin 3.967 (0.460)
thiothixene 2.546 (0.360)
troglitazone 3.329 (0.228)
valsartan 3.066 (1.031)
zaleplon 3.461 (0.501)
B.4 Extended Results — Trypsin inhibitors
@@
Appendix C Hyperparameter Settings
@@
GP-MoLFormer embedding dimensionality Float 768 See [ross2025gpmolformerfoundationmodelmolecular]